A sum of $500 is in the form of denominations of $5 and $10. If the total number of notes is 90, find the number of notes of each type.
Answer by TimothyLamb(4379)

x = number of 5 dollar notes
y = number of 10 dollar notes
---
x + y = 90
5x + 10y = 500
---
put the system of linear equations into standard form
---
x + y = 90
5x + 10y = 500
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = number of 5 dollar notes = 80
y = number of 10 dollar notes = 10
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---